Solve the third-order initial value problem below using the method of Laplace transforms. y! + 2y – 11y – 12y = - 48, y(
Answer #1

The given diffrential equation is y424 ally-12y = -48 960)=7 y 10) = 4,4 (0) = 8 Taking laplace transform on both side L[8%+232-13 -12] F13) -8782-48-80-148-8 +77 [88+28° -11 5 –12] F18) – 98-183 -10 = -48 : [ g 3 +28²-113-12] f 13 = -48 + (4 g* 3418)} = 44-46 484 - 4+ +(4+tuett (4+ na jest +14+ $) ett YH) = 44 - 49 6 + - 143 at +62 est TR2 Hence solve. od 0